OpenStudy (anonymous):

How can I use substitution to evaluate the integral from 0 to 3pi/8 of sin(2x-pi/4)dx?

zepdrix (zepdrix):

You don't need substitution :) Just a simple trick to remember. I'll explain... But if you insist on sub, we can do that also.

OpenStudy (anonymous):

The problem specifies substitution, but I would appreciate a short explanation on how to do it without as well, if you could :)

zepdrix (zepdrix):

\[\large\rm \frac{d}{dx}[-\cos(2x)]\]Do you understand how to do this derivative? :)

OpenStudy (anonymous):

Yes, with the chain rule, right? If I'm correct it would be -2sin2x.

zepdrix (zepdrix):

-cosine goes back to sine so yes, that's close :)

OpenStudy (anonymous):

Yes, I realized my mistake right away

zepdrix (zepdrix):

So if the derivative of -cos(2x) is 2sin(2x), then we expect that the integral of sin(2x) will be `something like` cos(2x) right? Maybe a negative, we're not sure exactly what is happening with the 2 as well. But that is a good "guess" at this point, right?

OpenStudy (anonymous):

Right

zepdrix (zepdrix):

Anyway, I'll cut to the point. Whenever you have these `linear coefficients`, differentiation gives you an extra coefficient on the outside due to chain rule. Think of integration doing the opposite. You simply end up dividing by that linear coefficient. Example:\[\large\rm \int\limits e^{2x}dx\quad=\quad \frac{1}{2}e^{2x}+C\]Differentiation would give us `multiplication of` 2 on the outside. Integration gives us `division by` 2 on the outside.

zepdrix (zepdrix):

Another example:\[\large\rm \int\limits \cos(7x+\pi)dx\quad=\quad \frac{1}{7}\sin(7x+\pi)\]Differentiation would normally give us an extra 7 on the outside, so integration must be taking away a 7.

OpenStudy (anonymous):

Ok, so that would mean the integral of sin(2x) would be -(cos(2x))/2, right?

zepdrix (zepdrix):

Ok great! We should probably do substitution to verify this process though, right? :) And since the instructions told us to. hehe

OpenStudy (anonymous):

Sorry if it is a little bit difficult to read, I'm not sure how to type it so that it will display it more clearly as an equation

OpenStudy (anonymous):

Yea, substitution would be great ha

zepdrix (zepdrix):

\[\large\rm \int\limits_0^{3\pi/8} \sin\left(\color{orangered}{2x-\frac{\pi}{4}}\right)\color{royalblue}{dx}\]

zepdrix (zepdrix):

\[\large\rm \color{orangered}{u=2x-\frac{\pi}{4}}\]We want to take the derivative of our substitution, and try to end up with the blue piece in our integral.

OpenStudy (anonymous):

So du would be x^2, right?

zepdrix (zepdrix):

Woops, differentiation :))

zepdrix (zepdrix):

x's get smaller when differentiating, ya?

OpenStudy (anonymous):

My bad, du = 2

zepdrix (zepdrix):

\[\large\rm du=2dx\]Ok that's almost exactly what we wanted. Maybe we can `solve` for dx here?

OpenStudy (anonymous):

So dx would be du/2?

zepdrix (zepdrix):

\[\large\rm \color{royalblue}{\frac{1}{2}du=dx}\]Ok great!

zepdrix (zepdrix):

\[\large\rm \int\limits\limits_0^{3\pi/8} \sin\left(\color{orangered}{2x-\frac{\pi}{4}}\right)\color{royalblue}{dx}\quad=\quad \int\limits\limits \sin\left(\color{orangered}{u}\right)\color{royalblue}{\frac{1}{2}du}\]

zepdrix (zepdrix):

Do you understand why I erased the limits?

OpenStudy (anonymous):

I think so. The way it is written without the limits now, I can take the indefinite integral first, then evaluate it later, right?

zepdrix (zepdrix):

Well the only reason I pointed that out is because those old limits `were values for x`. You can't plug them in for the variable u. So yes, you have a couple options: 1. Find new limits for u, integrate, evaluate using new limits. 2. Integrate, undo substitution, use original limits.

OpenStudy (anonymous):

I think I like using that second method better

zepdrix (zepdrix):

I hope my example didn't mess you up from earlier. We agreed that the antiderivative of sin(2x) was -(cos2x)/2. But let's try to match it with this problem a little better. sin(2x+stuff) we expect to give us -(cos(2x+stuff))/2 The inner function won't change, just a quick note there :)

OpenStudy (anonymous):

I think I'm following along alright so far., yes.

OpenStudy (anonymous):

I can factor out the 1/2 and put it in front of the integral, right?

zepdrix (zepdrix):

\[\large\rm =\frac12\int\limits \sin u~du\]Sounds good.

OpenStudy (anonymous):

Thanks for the help :)

zepdrix (zepdrix):

Got it from there? c:

OpenStudy (anonymous):

Yes, it makes sense from here, thanks

zepdrix (zepdrix):

cool